OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> 3.0° 6.0° -7.0° 16.0°

OpenStudy (amistre64):

\[\cos(\alpha)=\frac{u\cdot v}{|u||v|}\]

OpenStudy (anonymous):

okay?

OpenStudy (amistre64):

if you have no idea what that means, then youve apparently have either not reached that level, or you dont want to do the work .... what methods do you know of that you could apply?